A system of diatomic ideal gas is in an initial state such that the pressure is 69.0 kPa and the volume occupied by the gas is 6.00 L. The system then experiences a compression at constant temperature that raises the pressure to 165 kPa.

(a) Calculate the final volume occupied by the gas.
__L

(b) Calculate the work done by the gas in this process. (Include the sign of the value in your answer.)
__J

Please show all work!
Answer #1

let,
initial pressure P1=69kPa

final pressure P2=165kPa

initial volume v1=6L


a)

by using ideal gas equation,

P*V=n*R*T

==>

P1*v1=P2*v2

69*6=165*v2

==> v2=2.51L

final volume v2=2.51L


b)


in isothermal process,

workdone W=n*R*T*ln(V2/v1)

W=P1*V1*ln(4.1/1.3)

W=(69*10^3)*(6*10^-3)*ln(2.51/6)

W=-360.8 J